25. Today (1-0), an investor purchased a 20 year bond with a 5.00% coupon and a face value of $100,000 for $106,550. In six months (T 0.5) interest rates have decreased by 0.50% and the investor decides to sel the bond immediately after receiving the first coupon payment. What is the investors total gain (loss) on the bond? HINT: Total Gain (Loss)Price Change in Bond +Coupon A. ($6,548) B. ($6,048) C. $7,130 D. $7,602 E. $7,630 Assume all future cash flows (FCFF, FCFE, and Dividends) will grow at a constant sustainable growth rate (g) in perpetuity; note this growth rate is dependent on a companys return on equity and dividend payout policy. The companys cost of debt is 20% while the companys cost of equity is 30%; the company has an effective tax rate of 35%. Use the following information in the table below to help answer problems 26-27: FCFE Today (T 0) FCFF Today (T 0) Shareholder Equity Total Debt Total Assets Net Income Dividends Shares Outstanding in millions 800 750 400 600 1000 100 40 26. An activist investor who wants to purchase all the companys shares would be willing to pay approximately . Do not use the DDM. Use either FCFF or FCFE, whichever is appropriate. a. $165 per share b. $176 per share c. $213 per share d. $230 per share e. $245 per share 27. A private equity group who wants to purchase all of the companys assets would be willing to pay approximately. Do not use the DDM. Use either FCFF or FCFE, whichever is appropriate a. $9,583M b. $10,222M c. $10,518M d. $17,969M e. $19,167M
Answer #1

Question 25:

Fist we calculate the YTM at the time of sale

FV = 100,000

PV =106,550

PMT = 5% *100,000 = 5,000/2 = 2,500 (semi annually)

NPER = 20  years = 20 *2 = 40 semi annual periods

YTM of the the bond is calculated using =RATE(nper,pmt,pv,fv) *2=RATE(40,2500,-106550,100000)*2 = 4.49999 = 4.50%

YTM = 4.50%

Interest rates decrease by 0.50%. So new YTM or RATE = 4.00%

Now, we calculate the price using =PV(rate,nper,pmt,fv) where rate = 4.00%/2,nper =39,pmt = 2500 and fv =100000.

Price after six months at sale =PV(4%/2,39,2500,100000) = 113,451.29

So , Total Gain = 113,451.29 -106,550 + 2500 = 9,403

So, nearest answer choice is E

Note: We have answered one full question. Please note that only one MCQ can be answered at a time. Kindly post others separately for experts to answer